OpenStudy (anonymous):

Which of the following sets of numbers satisfy the conditions to be a "Pythagorean Triple?"

OpenStudy (anonymous):

3,6,7 5,12, 13 5,7,8 12,24,25 all of the above

OpenStudy (anonymous):

The pythagorean theorem states: \[a ^{2}+b ^{2}=c ^{2}\] so substituting in the numbers you have: 3^2 + 6^2 = 7^2 9+36=49 45=49 which is false so the first set does not work. See if you can try the other sets following the example.

OpenStudy (amistre64):

being a carpenter tells me 5-12-13

OpenStudy (anonymous):

so that is why you're good at Geometry.

OpenStudy (anonymous):

5^2 + 12^2 = 13^2 25 + 144 = 169 169 = 169 So it is the 5, 12, 13